OpenStudy (anonymous):

What is the standard form equation of the line passing through (–2, 8) with a slope of 2

OpenStudy (zzr0ck3r):

y = mx+b y = 2x+b what is b? use the point 8 = -2(2)+b b = 12 so y = 2x+12